# A primer on inference and prediction with epidemic renewal models and sequential Monte Carlo

### v1

@@ 4.1.1 Practical considerations
In addition to the choice of fixed lag $L$, there are a number of practical aspects to consider.
A greater number of particles $N$ will improves the accuracy of the posterior distribution approximation but increases computation time. The appropriate value of $N$ depends on the complexity of the model, how well the model fits the data, and the purpose of the bootstrap filter. In practice, we find that $N=100{,}000$ is sufficient (i.e., produces stable central estimates and credible intervals) for all models considered in this primer. This can be decreased (typically to around $N=1{,}000$) when the algorithm is used for likelihood estimation (section 4.2) or when the algorithm is used to find the marginal smoothing posterior distribution (section 5.1). The choice of $N$ for a specific model can be guided by comparing hidden-state estimates across different runs: if the estimates are stable (i.e., they do not noticeably change between runs), then $N$ is likely sufficiently large. Further validation can be performed by simulating data from the model and checking that the model recovers the simulated truth, or by fitting a similar model with a known analytical solution (such as EpiFilter) and comparing results to the known ground truth.
In algorithm 1, for simplicity, we use the state-space transition distribution as the proposal distribution. This is known to be suboptimal, particularly because it does not use information about the observed data $y_{t}$ when proposing new particles. More complex proposal distributions can be employed and the weights adjusted to account for the fact that the proposal distribution is not the true state-space transition distribution. One such example is the auxiliary particle filter 52, although we find this unnecessary for the models fit in this primer.
Due to fixed-lag resampling, estimates are usually robust to reasonable choices of the initial state distribution $P(X_{0})$. In the absence of clear prior information, it should allow for any plausible value of $X_{0}$.
@@ 4.2 Particle marginal Metropolis-Hastings
The log-likelihood estimator is given by the logarithm of the geometric mean of particle weights $\bar{w}_{t}$ from the bootstrap filter (algorithm 1) at each time step:
@@
where a Monte Carlo estimator of the one-step-ahead predictive likelihood is given by the average of the weights of the projected particles $\bar{w}_{t}$:
@@
$\approx\frac{1}{N}\sum_{i=1}^{N}P(y_{t}|\tilde{x}_{i,t},y_{1:t-1},\theta)$
@@
Algorithm 2 Particle marginal Metropolis Hastings (PMMH). The bootstrap filter $\mathcal{M}$ written here accepts proposed parameters $\theta^{\prime}$ and returns a matrix of weights. The data $y_{1:T}$ and other bootstrap filter options are assumed to be included in $\mathcal{M}$.
1:Input: Number of parameter samples $M$, bootstrap filter $\mathcal{M}$, parameter prior distribution $P(\theta)$, parameter proposal distribution $P(\theta^{\prime}|\theta)$.
@@
4:     $\theta^{\prime}\sim P(\theta^{\prime}|\theta_{j-1})$triangleright Sample proposed parameters
5:     $W\leftarrow\mathcal{M}(\theta^{\prime})$triangleright Run the bootstrap filter (algorithm 1) at $\theta^{\prime}$
6:     ^$\hat{\ell}(\theta^{\prime})=\leftarrow\frac{1}{T}\sum_{t=1}^{T}\log\bar{W}_{\cdot,t}$triangleright Calculate log-likelihood estimate (equation 2)
7:     $\alpha\leftarrow\min\left(\frac{\exp\hat{\ell}(\theta^{\prime})}{\exp\hat{\ell}(\theta_{j-1})}\frac{P(\theta^{\prime})}{P(\theta_{j-1})}\frac{P(\theta_{j-1}|\theta^{\prime})}{P(\theta^{\prime}|\theta_{j-1})},1\right)$triangleright Calculate acceptance probability
@@ 4.2.1 Practical considerations
To enable the calculation of convergence diagnostics, multiple chains (we find 4 works well) of the PMMH algorithm should be run. We also encourage the standard practice of discarding an initial portion of each chain as a burn-in period, ensuring the retained samples are not influenced by initial values outside the stationary distribution of the Markov chain. Diagnostics used in the examples below are the Gelman-Rubin statistic $\hat{R}$ 55 and effective sample size (ESS) 54, which are reported by standard Markov chain Monte Carlo (MCMC) analysis software. We use the MCMCChains.jl package 56 in our examples.
The standard Metropolis-Hastings algorithm uses exact evaluations of the model log-likelihood $\ell(\theta)$, whereas algorithm 1 admits a stochastic estimator ^$\hat{\ell}(\theta)$. The standard deviation of this estimator is a function of the number of particles $N$ used in the bootstrap filter. It has been shown that the optimal standard deviation of ^$\hat{\ell}(\theta)$ is 1.2-1.3 43, 57, which can be used to guide the choice of $N$. For all examples in this primer, we use $N=1,000$ particles when estimating model likelihoods. We also note that, while larger standard deviations of ^$\hat{\ell}(\theta)$ result in slower convergence, model outputs are still valid.
The efficiency of the PMMH algorithm depends on the parameter proposal distribution $P(\theta^{\prime}|\theta)$. We find that the heuristic multivariate normal proposal distribution with covariance matrix $\Sigma=(2.38^{2}/d)\hat{\Sigma}$, where $\hat{\Sigma}$ is the sample covariance matrix of previous samples and $d$ is the number of parameters being fit, generally performs well 58.
@@ 5.2.1 Predictive posterior distribution of observed data
Various predictive posterior distributions can be targeted.
For example, the one-step-ahead predictive posterior distribution $P(Y_{t}|y_{1:t-1},\theta)$ is targeted by sampling from the observation distribution given the projected particles during the bootstrap filter. That is, sampling $y_{t}^{(i)}\sim P(Y_{t}|\tilde{x}_{1:t}^{(i)},\theta)$, where $\tilde{x}_{1:t}^{(i)}$ are the projected particles as defined in algorithm 1.
Alternatively, the smoothing posterior predictive distribution $P(Y_{t}|y_{1:T},\theta)$ is targeted by sampling from the observation distribution given the particles at time step $t$, that is $y_{t}^{(i)}\sim P(Y_{t}|x_{1:t}^{(i)},\theta)$, and storing and resampling these samples alongside the hidden states.
In both cases, the parameter vector $\theta$ can be marginalised out by including and storing the relevant samples in algorithm 3.
@@ 5.2.2 Missing data
Missing data can be handled in a variety of ways depending on the type of missingness (missing at random, missing not at random, etc.59). Where data are missing by design (for example, wastewater sampling data may only be collected on certain days 33), the observation distribution becomes $P(y_{t}\text{ is missing}|X_{1:t},y_{1:t-1},\theta)=1$, irrespective of the values of the hidden states. In practice, we also skip the resampling step in the bootstrap filter on such days, as all hidden states are equally likely. More complex models, where the relationship between hidden states and missingness is explicitly modelled, are also possible. Prediction of missing data can be made by sampling from the observation model, conditional on the value of the hidden states (another type of predictive posterior distribution).
@@ 5.2.4 Probability of elimination
Elimination of an infectious disease is generally defined as “the reduction to zero incidence of a certain pathogen in a given area” 60. For the purposes of this primer, we say that elimination has occurred at time $t$ if no new infections occur within $[t,t+28]$, an appropriate window for the SARS-CoV-2 examples we consider, although a variety of other mathematical statements matching this definition are possible 34, 9. This turns the problem of estimating the probability of elimination into a projection problem: we project the number of new infections in the next 28 days, and the proportion of particle trajectories that contain zero new infections provides an estimate of the probability of elimination. We demonstrate this in example (2) below.
@@ 5.3 Model evaluation and selection
Model evaluation and selection are crucial components of any modelling exercise. We outline three metrics: root mean square error (RMSE) of the posterior predictive distribution, the coverage of posterior predictive credible intervals, and the continuous ranked probability score (CRPS) 61 of the posterior predictive distribution.
@@
In practice, we use a particle approximation to the expectation-definition of the CRPS 62:
@@
where $Y_{t}^{(j)}$ are samples from the posterior predictive distribution at time $t$ and $\tilde{Y}^{(j)}$ are the samples sorted in ascending order. Using sorted samples reduces the computational complexity of the CRPS from $\mathcal{O}(N^{2})$ (line 2) to $\mathcal{O}(N\log N)$ (line 3) 62. The CRPS is particularly useful in the model selection process, as it provides a principled trade-off between precision (how narrow the credible intervals are) and calibration.
@@ 6 Example: COVID-19 in Aotearoa New Zealand
We demonstrate these methods by fitting three models to national data from the COVID-19 pandemic in Aotearoa New Zealand 63. Two time periods are considered: 26 February 2020 - 4 June 2020 and 1 April 2024 - 9 July 2024. The former is characterised by a single epidemic wave with high-quality case reporting and a large proportion of imported cases, during which elimination of transmission was the primary public health policy aim. The latter is characterised by widespread transmission with a clear day-of-week effect and high levels of reporting noise and bias, with modelling primarily used to inform health-service resource allocation.
Figure 5: Reported cases from the COVID-19 pandemic in New Zealand, separated by locally-acquired and imported infections, for two time-periods: (A) 26 February 2020 - 4 June 2020 and (B) 1 April 2024 - 9 July 2024 63. The first period is the first wave of COVID-19 in New Zealand, characterised by a single epidemic wave which ended in temporary elimination of local transmission. The second period is characterised by widespread transmission with much higher numbers of reported cases.
@@ 6.1 Model 1: Simple example
As a first example, we implement a model very similar to EpiFilter 5, assuming that $R_{t}$ follows a log-normal random walk (the state-space transition model):
@@ 6.2 Model 2: Reporting noise, imported cases, and elimination probabilities
During this period, elimination of SARS-CoV-2 was an explicit public-health policy goal in New Zealand. Models explicitly estimating the probability of elimination were of key interest to policymakers 64, particularly as repeated days of zero cases were observed. We demonstrate how these methods can be used to estimate the probability of elimination, defined here as the probability that the true infection incidence $I_{t}$ is zero for the next 28 days (i.e. $I_{t:t+28}=0$), while simultaneously accounting for reporting noise and imported cases.
@@
Using PMMH (algorithm 2) we estimate a posterior mean and 95% credible interval of $\sigma$ of 0.18 (95% Cr.I. 0.11, 0.29) and $\phi$ of 0.018 (95% Cr.I. 0.0007, 0.067). The decreased estimate of $\sigma$ (compared to model 1) reflects the re-attribution of some noise in the data from the epidemic process to the observation process. The estimate of $\phi$ is small, suggesting that the observation process may be adequately modelled by a Poisson distribution, instead of the negative binomial distribution we have assumed (this could be tested using the model selection metrics outlined in section 5.3). Convergence was obtained in 800 iterations, taking approximately 50 seconds on a 2021 M1 MacBook Pro.
Daily posterior means and 95% credible intervals of $R_{t}$ are reported in figure 7-A. Despite being fit to data from the same outbreak as model 1, the estimates of $R_{t}$ are very different. By halting the algorithm on 5 April 2020, we estimate that $R_{t}$ peaked at 1.7 (95% Cr.I. 1.2, 2.4) on 23 March 2020 (95% Cr.I. 17 March, 26 March), much lower and somewhat later than estimates from the simple model. This decrease is primarily a result of distinguishing locally-acquired from imported cases, although the re-attribution of noise from the epidemic process to the observation process also plays a role.
@@
The probability of elimination increases steadily from early May and is estimated to reach 84.3% on the final time step 4 June 2020. Conditioning on no new imported cases is equivalent to assuming that any imported infections cause no local transmission, a reasonable assumption given the strict border controls in place at the time.

### v2

@@ 4.1.1 Practical considerations
The choice of fixed lag $L$ depends on the modelled maximum serial interval (or generation time) $u_{max}$. In many cases, such as in the COVID-19 examples considered, $u_{max}$ (and thus $L$) can be chosen to be sufficiently small to ensure algorithmic efficiency without introducing noticeable bias. Some diseases, such as tuberculosis 60, feature serial intervals and generation times that are measured in weeks, months, or even years. While this poses a computational problem when using a daily discretisation (requiring a very large $L$), the longer serial intervals/generation times allow for the pathogen to be modelled on a weekly (or even monthly) discretisation, thus $L$ can be kept small. The impact of $L$ can be checked by refitting the model multiple times at different values of $L$ and comparing the outputs. This is demonstrated online, where we find that reasonable choices of $L$ have very little impact on an example model.
A greater number of particles $N$ improves the accuracy of the posterior distribution approximation but increases computation time. The appropriate value of $N$ depends on the complexity of the model, how well the model fits the data, and the purpose of the bootstrap filter, and thus is difficult to determine a priori. In practice, we find that $N=100{,}000$ is sufficient (i.e., produces stable central estimates and credible intervals) for all models considered in this primer. This can be decreased (typically to around $N=1{,}000$) when the algorithm is used for likelihood estimation (see Section 4.2 for further discussion) or when the algorithm is used to find the marginal smoothing posterior distribution (Section 5.1).
The choice of $N$ for a specific model can be guided by comparing hidden-state estimates across multiple runs: if the estimates are stable (i.e., they do not noticeably change between runs), then $N$ is likely sufficiently large. If a programmatic heuristic is required, the particle weights (line 5 of Algorithm 1) can be used to calculate the effective sample size61, which is inversely proportional to the variance of the estimate of the posterior mean and credible interval width. The calculation of effective sample size is demonstrated on an example model online. Further validation can be performed by simulating data from the model and checking that the model recovers the simulated truth, or by fitting a similar model with a known analytical solution (such as EpiFilter) and comparing results to the known ground truth. Theoretical convergence results 62 developed for Markovian state-space models can be applied by augmenting past states into $X_{t}$ and considering the model on the expanded state space.
In Algorithm 1, for simplicity, we use the state-space transition distribution as the proposal distribution. This is known to be suboptimal, particularly because it does not use information about the observed data $y_{t}$ when proposing new particles. More complex proposal distributions can be employed and the weights adjusted to account for the fact that the proposal distribution is not the true state-space transition distribution. One such example is the auxiliary particle filter 63, although we find this unnecessary for the models fit in this primer.
Due to fixed-lag resampling, estimates are usually robust to reasonable choices of the initial state distribution $P(X_{0})$. In the absence of clear prior information, the initial state distribution should allow for any plausible value of $X_{0}$. The effect of the initial state distribution can be examined through sensitivity analysis.
@@ 4.2 Particle marginal Metropolis-Hastings
The log-likelihood estimator is given by the logarithm of the geometric mean of the unnormalised particle weights $\bar{w}_{t}$ from the bootstrap filter (Algorithm 1) at each time step 38, 66:
@@
where a Monte Carlo estimator of the one-step-ahead predictive likelihood is given by the average of the unnormalised weights of the projected particles, denoted $\bar{w}_{t}$:
@@
$\approx\frac{1}{N}\sum_{i=1}^{N}P\left(y_{t}|\tilde{x}_{t}^{(i)},y_{1:t-1},\theta\right)$
@@
Algorithm 2 Particle marginal Metropolis Hastings (PMMH). The bootstrap filter $\mathcal{M}$ written here accepts proposed parameters $\theta^{\prime}$ and returns unnormalised weights $\{w_{t}^{(i)}\}$ for $i=1,\ldots,N$ and $t=1,\ldots,T$. The average of all weights at time $t$ is denoted $\bar{w}_{t}$. The data $y_{1:T}$ and other bootstrap filter options are assumed to be included in $\mathcal{M}$.
1:Input: Number of parameter samples $M$, bootstrap filter $\mathcal{M}$, parameter prior distribution $P(\theta)$, parameter proposal distribution $q(\theta^{\prime}|\theta)$.
@@
4:     $\theta^{\prime}\sim q(\theta^{\prime}|\theta_{j-1})$triangleright Sample proposed parameters
5:     $\{w_{t}^{(i)}\}_{i=1,\ldots,N,\ t=1,\ldots,T}\leftarrow\mathcal{M}(\theta^{\prime})$triangleright Run the bootstrap filter (Algorithm 1) at $\theta^{\prime}$
6:     ^$\hat{\ell}(\theta^{\prime})=\leftarrow\frac{1}{T}\sum_{t=1}^{T}\log\bar{w}_{t}$triangleright Calculate log-likelihood estimate (equation 2)
7:     $\alpha\leftarrow\min\left(\frac{\exp\hat{\ell}(\theta^{\prime})}{\exp\hat{\ell}(\theta_{j-1})}\frac{P(\theta^{\prime})}{P(\theta_{j-1})}\frac{q(\theta_{j-1}|\theta^{\prime})}{q(\theta^{\prime}|\theta_{j-1})},1\right)$triangleright Calculate acceptance probability
@@ 4.2.1 Practical considerations
To enable the calculation of convergence diagnostics, multiple chains (we find 4 works well) of the PMMH algorithm should be run. We also encourage the standard practice of discarding an initial portion of each chain as a burn-in period, ensuring the retained samples are not influenced by initial values outside the stationary distribution of the Markov chain. Diagnostics used in the examples below are the Gelman-Rubin statistic $\hat{R}$ 67 and effective sample size (ESS) 65, which are reported by standard Markov chain Monte Carlo (MCMC) analysis software. We use the MCMCChains.jl package 68 in our examples.
The standard Metropolis-Hastings algorithm uses exact evaluations of the model log-likelihood $\ell(\theta)$, whereas Algorithm 1 admits a stochastic estimator ^$\hat{\ell}(\theta)$. The standard deviation of this estimator is a function of the number of particles $N$ used in the bootstrap filter. It has been shown that the optimal standard deviation of ^$\hat{\ell}(\theta)$ is 1.2-1.3 45, 69, which can be used to guide the choice of $N$. The standard deviation of the log-likelihood estimate can itself be estimated by refitting the model multiple times, we demonstrate this online. For all examples in this primer, we use $N=1,000$ particles when estimating model likelihoods. We also note that, while larger standard deviations of ^$\hat{\ell}(\theta)$ result in slower convergence, model outputs are still valid.
The efficiency of the PMMH algorithm depends on the parameter proposal distribution $q(\theta^{\prime}|\theta)$. We find that the heuristic multivariate normal proposal distribution with covariance matrix $\Sigma=(2.38^{2}/d)\hat{\Sigma}$, where $\hat{\Sigma}$ is the sample covariance matrix of previous samples and $d$ is the number of parameters being fit, generally performs well 70.
@@ 5.2.1 Predictive posterior distribution of observed data
Various predictive posterior distributions can be targeted. For example, the one-step-ahead predictive posterior distribution $P(Y_{t}|y_{1:t-1},\theta)$ can be targeted within the bootstrap filter by sampling from the observation distribution given the projected particles. That is, sampling $y_{t}^{(i)}\sim P(Y_{t}|\tilde{x}_{1:t}^{(i)},\theta)$, where $\tilde{x}_{1:t}^{(i)}$ are the projected particles as defined on line 4 of Algorithm 1.
Alternatively, the smoothing posterior predictive distribution $P(Y_{t}|y_{1:T},\theta)$ can be targeted within the bootstrap filter by sampling from the observation distribution given the projected particles at each time step $t$ as above (setting $\tilde{y}_{t}^{(i)}\sim P(Y_{t}|\tilde{x}_{1:t}^{(i)},\theta)$). Defining $\tilde{y}_{t-L:t}^{(i)}$ as the aggregation of the newly sampled $\tilde{y}_{t}^{(i)}$ and previously stored $y_{t-L:t-1}^{(i)}$, and resampling $y_{t-L:t}^{(i)}$ from $\tilde{y}_{t-L:t}^{(i)}$ alongside $x_{t-L:t}^{(i)}$ in step 6 of the Algorithm 1 results in a set of particles $\{y_{t}^{(i)}\}$ that can be viewed as unweighted samples from the smoothing posterior predictive distribution. This can be clearly seen if we consider the predictive variable as part of an extended hidden state space. To ensure consistency between the weights calculated in step 5 of Algorithm 1 and the predictive particles, it is important to use the same indices when resampling both $x_{t}^{(i)}$ and $y_{t}^{(i)}$. Computationally this is equivalent to sampling some indices from $\text{Multinomial}\left(\{1,2,\ldots,N\},\left\{\frac{w_{t}^{(j)}}{\sum_{k=1}^{N}w_{t}^{(k)}}\right\}_{j=1}^{N}\right)$, and then using these to resample both the hidden states and predictive variable. The parameter vector $\theta$ can be marginalised out by including and storing the relevant samples in Algorithm 3.
@@ 5.2.2 Missing data
Missing data can be handled in a variety of ways depending on the type of missingness (missing at random, missing not at random, etc.71). Where data are missing by design (for example, wastewater sampling data may only be collected on certain days 23), the observation distribution becomes $P(y_{t}\text{ is missing}|X_{1:t},y_{1:t-1},\theta)=1$, irrespective of the values of the hidden states. In practice, we also skip the resampling step in the bootstrap filter on such days, as all hidden states are equally likely. More complex models, where the relationship between hidden states and missingness is explicitly modelled, are also possible. Prediction of missing data can be made by sampling from the observation model, conditional on the value of the hidden states (another type of predictive posterior distribution).
@@ 5.2.4 Probability of elimination
Elimination of an infectious disease is generally defined as “the reduction to zero incidence of a certain pathogen in a given area” 72. For the purposes of this primer, we say that elimination has occurred at time $t$ if no new infections occur within $[t,t+28]$, an appropriate window for the SARS-CoV-2 examples we consider, although a variety of other mathematical statements matching this definition are possible 36, 11. This turns the problem of estimating the probability of elimination into a projection problem: we project the number of new infections in the next 28 days, and the proportion of particle trajectories that contain zero new infections provides an estimate of the probability of elimination. We demonstrate this in example (2) below.
@@ 5.3 Model evaluation and selection
Model evaluation and selection are crucial components of any modelling exercise. We outline three metrics: root mean square error (RMSE) of the posterior predictive distribution, the coverage of posterior predictive credible intervals, and the continuous ranked probability score (CRPS) 73 of the posterior predictive distribution.
@@
In practice, we use a particle approximation to the expectation-definition of the CRPS 74:
@@
where $Y_{t}^{(j)}$ are samples from the posterior predictive distribution at time $t$ and $\tilde{Y}^{(j)}$ are the samples sorted in ascending order. Using sorted samples reduces the computational complexity of the CRPS from $\mathcal{O}(N^{2})$ (line 2) to $\mathcal{O}(N\log N)$ (line 3) 74. The CRPS is particularly useful in the model selection process, as it provides a principled trade-off between precision (how narrow the credible intervals are) and calibration.
@@ 6 Example: COVID-19 in Aotearoa New Zealand
We demonstrate these methods by fitting three models to national data from the COVID-19 pandemic in Aotearoa New Zealand 75. Two time periods are considered: 26 February 2020 - 4 June 2020 and 1 April 2024 - 9 July 2024. The former is characterised by a single epidemic wave with high-quality case reporting and a large proportion of imported cases, during which elimination of transmission was the primary public health policy aim. The latter is characterised by widespread transmission with a clear day-of-week effect and high levels of reporting noise and bias, with modelling primarily used to inform health-service resource allocation.
Figure 5: Reported cases from the COVID-19 pandemic in New Zealand, separated by locally-acquired and imported infections, for two time-periods: (A) 26 February 2020 - 4 June 2020 and (B) 1 April 2024 - 9 July 2024 75. The first period is the first wave of COVID-19 in New Zealand, characterised by a single epidemic wave which ended in temporary elimination of local transmission. The second period is characterised by widespread transmission with much higher numbers of reported cases.
@@ 6.1 Model 1: Simple example
As a first example, we implement a model very similar to EpiFilter 7, assuming that $R_{t}$ follows a log-normal random walk (the state-space transition model):
@@ 6.2 Model 2: Reporting noise, imported cases, and elimination probabilities
During this period, elimination of SARS-CoV-2 was an explicit public-health policy goal in New Zealand. Models explicitly estimating the probability of elimination were of key interest to policymakers 76, particularly as repeated days of zero cases were observed. We demonstrate how these methods can be used to estimate the probability of elimination, defined here as the probability that the true infection incidence $I_{t}$ is zero for the next 28 days (i.e. $I_{t:t+28}=0$), while simultaneously accounting for reporting noise and imported cases.
@@
Using PMMH (Algorithm 2) we estimate a posterior mean and 95% credible interval of $\sigma$ of 0.16 (95% Cr.I. 0.09, 0.25) and $\phi$ of 0.018 (95% Cr.I. 0.0005, 0.065). The decreased estimate of $\sigma$ (compared to model 1) reflects the re-attribution of some noise in the data from the epidemic process to the observation process. The estimate of $\phi$ is small, suggesting that the observation process may be adequately modelled by a Poisson distribution, instead of the negative binomial distribution we have assumed (this could be tested using the model selection metrics outlined in Section 5.3). Convergence was obtained in 800 iterations, taking approximately 50 seconds on a 2021 M1 MacBook Pro.
Daily posterior means and 95% credible intervals of $R_{t}$ are reported in Figure 7-A. Despite being fit to data from the same outbreak as model 1, the estimates of $R_{t}$ are very different. By halting the algorithm on 5 April 2020, we estimate that $R_{t}$ peaked at 1.6 (95% Cr.I. 1.2, 2.2) on 23 March 2020 (95% Cr.I. 6 March, 27 March), much lower and somewhat later than estimates from the simple model. This decrease is primarily a result of distinguishing locally-acquired from imported cases, although the re-attribution of noise from the epidemic process to the observation process also plays a role.
@@
The probability of elimination increases steadily from early May and is estimated to reach 91.8% on the final time step 4 June 2020. Conditioning on no new imported cases is equivalent to assuming that any imported infections cause no local transmission, a reasonable assumption given the strict border controls in place at the time.